OpenStudy (anonymous):

PLEASE HELP ME Where does the graph of y=2x^2 - 9x -5 cross the x-axis?

OpenStudy (anonymous):

crosses the \(x\) axis where \[2x^2-9x-5=0\] can you solve it?

OpenStudy (anonymous):

not too bad because this one factors as \((x-5)(2x+1)=0\)

OpenStudy (anonymous):

@satellite73 i need like two points

OpenStudy (anonymous):

yes you do

OpenStudy (anonymous):

solve \(x-5=0\) and solve \(2x+1=0\)

OpenStudy (anonymous):

These are the answer choices -1/2 ,0 and 5,0 1/2,0 and -5,0 -5/2,0 and 1,0 5/2,0 and -1,0

OpenStudy (anonymous):

Ohhhhh ok

OpenStudy (anonymous):

if it is on the \(x\) axis the second coordinate is surely 0 question is what is the first coordinate i am sure you can solve \(x-5=0\) right?

OpenStudy (anonymous):

Ok so one of the xs is 5

OpenStudy (anonymous):

yes \((5,0)\) is where it crosses \(x\) axis also \((-\frac{1}{2},0)\)

OpenStudy (anonymous):

i think it might have been confusing because they wanted a coordinate rather than just the \(x\) value

OpenStudy (anonymous):

Thank u so much @satellite73